Q. Find the dimension of the subspace of R4 that is span of the vectors
(█(1¦(-1)@0@1)), (█(2¦1@1@1)),(█(0¦0@0@0)),(█(1¦1@-2@-5))
1
Expert's answer
2019-07-17T10:15:00-0400
"A=\begin{pmatrix}
1 & -1 & 0 & 1\\
2 & 1 & 1 & 1\\
0 & 0 & 0 & 0\\
1 & 1 & -2 & -5
\end{pmatrix}"

2 line minus 1 line and 4 line plus 1 line multiplied by 5:

"A \sim \begin{pmatrix}
1 & -1 & 0 & 1\\
1 & 2 & 1 & 0\\
0 & 0 & 0 & 0\\
6 & -4 & -2 & 0
\end{pmatrix}"

4 line plus 2 line multiplied by 2:


"A \sim \begin{pmatrix}
1 & -1 & 0 & 1\\
1 & 2 & 1 & 0\\
0 & 0 & 0 & 0\\
8 & 0 & 0 & 0
\end{pmatrix}"

Hence, we have three linearly independent vectors "\begin{pmatrix} 1\\1\\0\\8 \end{pmatrix}, \begin{pmatrix} 0\\1\\0\\0 \end{pmatrix},\begin{pmatrix} 1\\0\\0\\0 \end{pmatrix}," and


"\operatorname{rank}A=3"

So, we have 3-dimension subspace of "\mathbb{R}^4"